OpenStudy (anonymous):

Suppose that polynomial functions f(x) and g(x) are each a difference of square terms. What can you say about the roots of f(x)*g(x)?

OpenStudy (anonymous):

what does Suppose that polynomial functions f(x) and g(x) are each a difference of square terms mean? \[f(x)=x^2-9\] \[f(x)=(x+2)^2-(x-3)^2\]

OpenStudy (anonymous):

I don't know. That is what the questions says and that is why I can't figure it out.